OpenStudy (anonymous):

Find x 10^log(x^2 + 2x) = 15

OpenStudy (anonymous):

10^log(x^2+2x) is the same as x^2+2x therefore, x^2+2x=15 x^2+2x-15=0 use quadratic formula: x= (-2 + sqrt(4+60))/2 or x= (-2 - sqrt(4+60))/2 x= (-2 + 8)/2 = 3 or x=(-2-8)/2= -5 x= 3 or -5

OpenStudy (anonymous):

if A=10^B then log(A)=B therefore A=10^log(A) that's where I got it :)